h/2 + h/5= 14 I can't find how to simplify the equition
Answer by RAY100(1637)

h/2 +h/5 = 14
multiply by 10 (LCD)
5h +2h = 140
7h =140
divide by 7
h=20
check
20/2+20/5=14, 10+4=14 ok